OpenStudy (anonymous):

Help please! If m<2=40, then m<3+m<4=

ganeshie8 (ganeshie8):

sure :) u got a diagram attached wid this question ?

ganeshie8 (ganeshie8):

you may take a screenshot and attach the pic if its possible..

ganeshie8 (ganeshie8):

oops ! u did that already :)

OpenStudy (anonymous):

Nvm whoops

ganeshie8 (ganeshie8):

first step : m<2 and m<5 are vertical angles, so they are equal : m<5 = 40

ganeshie8 (ganeshie8):

next, see that m<3 = 90, and m<4 = 90-40 = 50

ganeshie8 (ganeshie8):

so, m<3 + m<4 = 90+50 = 140

ganeshie8 (ganeshie8):

see if that makes more or less sense

ganeshie8 (ganeshie8):

Another quick way is to see that 2,3,4 form a straight line :- m<2 + m<3 + m<4 = 180 m<3 + m<4 = 180 - m<2 = 180 - 40 = 140

ganeshie8 (ganeshie8):

both methods give u same answer. pick which ever u feel easy to digest.. :)

OpenStudy (anonymous):

Why do you subtract 90-40? (new to this) is that just basic protocol for this type of problem?

ganeshie8 (ganeshie8):

thats a very good question !

ganeshie8 (ganeshie8):

i did 90-40, cuz : m<5 + m<4 = 90

ganeshie8 (ganeshie8):

next u going to ask, why is that true ?

ganeshie8 (ganeshie8):

thats true, cuz from the diagram we see that there is a 'little square' box at the angle 6. that 'little square' box means 90 degrees.

ganeshie8 (ganeshie8):

that 'little square' makes four right angles. right angle = 90 degrees.

ganeshie8 (ganeshie8):

yah sure :) takes time... but these are easy once u get some practice :)

ganeshie8 (ganeshie8):

first two are right last one is not right

ganeshie8 (ganeshie8):

lets see, wheteher <1 + <5 gives us 90 or not

ganeshie8 (ganeshie8):

<1 + <5 but <5 = <2 as they are vertical angles <1 + <2

ganeshie8 (ganeshie8):

yes, now look at diagram. <1 + <2 ... do they make 90 ?

ganeshie8 (ganeshie8):

the 'little square' box by 6 makes all four angles 90.

ganeshie8 (ganeshie8):

so, the 3rd statement is indeed True

OpenStudy (anonymous):

I see! Thank you so much, I actually got practice out of my practice worksheet. I appreciate you taking the time to help me

ganeshie8 (ganeshie8):

good :) np... you're wlc :)